### v3

# Scaling Cross-Environment Failure Reasoning Data for Vision-Language Robotic Manipulation

Abstract
Robust robotic manipulation requires reliable failure detection and recovery. Although recent Vision-Language Models (VLMs) show promise in robot failure detection, their generalization is severely limited by the scarcity and narrow coverage of failure data. To address this bottleneck, we propose an automatic framework for generating diverse robotic planning and execution failures across both simulated and real-world environments. Our approach perturbs successful manipulation trajectories to synthesize failures that reflect realistic failure distributions, and leverages VLMs to produce structured step-by-step reasoning traces. This yields FailCoT, a large-scale failure reasoning dataset built upon the RLBench simulator and the BridgeDataV2 real-robot dataset. Using FailCoT, we train Guardian, a multi-view reasoning VLM for unified planning and execution verification. Guardian achieves state-of-the-art performance on three unseen real-world benchmarks: RoboFail, RoboVQA, and our newly introduced UR5-Fail. When integrated with a state-of-the-art LLM-based manipulation policy, it consistently boosts task success rates in both simulation and real-world deployment. These results demonstrate that scaling high-quality failure reasoning data is critical for improving generalization in robotic failure detection. Code, Data, and Models available at https://www.di.ens.fr/willow/research/guardian/.
I Introduction
Recent advances in Large Language Models (LLMs) [mistralsmall, grattafiori2024llama3herdmodels] and Vision-Language Models (VLMs) [zhu2025internvl3exploringadvancedtraining] have significantly improved vision-language robotic manipulation. Nevertheless, existing models remain vulnerable to diverse failures [sinha2023systemlevelviewoutofdistributiondata, Kawaharazuka_2024, kroemer2020reviewrobotlearningmanipulation] such as incorrect task decomposition, object confusion, or unstable grasps, which compound over long horizons and degrade real-world reliability. As a result, automatic failure detection and recovery has received growing research attention [liu2023reflect, chen2024automatingrobotfailurerecovery, duan2025aha, agia2024unpackingfailuremodesgenerative, etukuru2024robotutilitymodelsgeneral, ifailsense2026, zeng2025vifailback].
Leveraging their strong generalization ability, LLMs and VLMs have been increasingly explored for failure detection. Some methods [liu2023reflect, etukuru2024robotutilitymodelsgeneral, duan2024manipulateanythingautomatingrealworldrobots], directly prompt pretrained foundation models to detect failures, optionally enhanced with chain-of-thought (CoT) reasoning [agia2024unpackingfailuremodesgenerative, nvidia2026cosmosreason2, cot2022] or multi-agent code generation [zhou2024code]. While promising, these approaches suffer from a large domain gap: robotic observations differ substantially from web-scale pretraining data, and accurate failure detection requires fine-grained, embodied reasoning beyond generic visual understanding. Therefore, recent work [duan2025aha, ifailsense2026, robofac2025] has shifted toward fine-tuning VLMs on robot failure datasets to better bridge the gap.
A fundamental bottleneck, however, is the scarcity of large-scale, high-quality failure data. Most robot learning datasets predominantly contain successful demonstrations [khazatsky2025droidlargescaleinthewildrobot, embodimentcollaboration2024openxembodimentroboticlearning, pumacay2024colosseumbenchmarkevaluatinggeneralization], providing limited failure examples. Collecting failures by rolling out policies is time-consuming and potentially unsafe, while manual curation [chen2024automatingrobotfailurerecovery, bu2025agibot] is labor-intensive and typically lacks diversity. Several prior approaches [duan2025aha, agia2024unpackingfailuremodesgenerative, dai2024racer] rely on simulated failure examples, but these suffer from sim-to-real gap [simtorealgapzhao2020], and provide limited coverage of both low-level execution errors and high-level planning failures [ifailsense2026].
To address these limitations, we propose an automatic failure generation framework that synthesizes diverse planning and execution failures across simulated and real-world environments. Starting from successful demonstrations, we procedurally perturb task plans and subtask executions to create realistic failures, augmenting each example with structured step-by-step reasoning traces. This enables the construction of FailCoT, a large-scale failure reasoning dataset containing over 30K training examples. It includes RLBench-Fail built using the RLBench simulator [james2019rlbenchrobotlearningbenchmark] and BridgeDataV2-Fail derived from the BridgeDataV2 real-robot dataset [walke2024bridgedatav2datasetrobot], see Figure 1. FailCoT provides balanced success and failure samples, multi-view visual observations, and explicit CoT supervision for plan and subtask-level verification.
Building on FailCoT, we develop Guardian, a multi-view reasoning VLM fine-tuned for unified planning and execution failure detection. Guardian formulates verification as a visual question answering problem: conditioned on task instructions, proposed plans or subtasks, and multi-view observations, it produces explicit reasoning traces to enhance predicting failures. To further support realistic evaluation, we introduce a new real-robot benchmark UR5-Fail constructed using the same failure generation framework. Guardian achieves state-of-the-art performance on three unseen real-world failure benchmarks, namely RoboFail [liu2023reflect], RoboVQA [sermanet2024] and UR5-Fail. When integrated as a plug-and-play verification module into a LLM-based manipulation system, Guardian improves task success in both simulation and real-robot experiments. Extensive ablations demonstrate the benefits of scaling structured, cross-environment failure reasoning data.
In summary, our contributions are three-fold:
  • We propose an automatic cross-environment failure synthesis framework that generates diverse planning and execution errors with structured reasoning supervision, resulting in the large-scale robot failure dataset FailCoT.
  • We develop Guardian, a multi-view reasoning VLM fine-tuned on the FailCoT dataset for unified planning and execution failure detection.
  • We show that scaling structured failure reasoning data yields state-of-the-art detection performance and improves task success when deployed as a plug-and-play verifier.
We will release datasets, code, and models.
Figure 1: Failure Data Generation Framework. We construct failure cases both online in simulation (RLBench), and offline on the real-world dataset (BridgeDataV2). For each positive example, given its correct plan and successful trajectory, we generate a corresponding incorrect plan and unsuccessful trajectory.
II Related Work
Vision-Language Robotic Manipulation. Recent advances in foundation models [zhu2025internvl3exploringadvancedtraining, openai2024gpt4ocard] have significantly improved vision–language robotic manipulation. End-to-end vision–language–action (VLA) policies such as Gr00T [groot2025] and $\pi_{0}$ [black2024pi] directly predict action sequences from 2D images and task instructions. To enhance spatial reasoning, 3D-based VLAs have further been proposed [garcia2025generalizablevisionlanguageroboticmanipulation, goyal2023rvtroboticviewtransformer]. Notably, 3D-LOTUS++ [garcia2025generalizablevisionlanguageroboticmanipulation] achieves state-of-the-art performance on challenging generalizable manipulation tasks through a modular design that combines LLM-based task planning, visual grounding modules, and 3D-based execution policies. Despite this progress, robust robotic manipulation remains challenging, as planning mistakes and execution errors accumulate over long horizons [wu2025robomind]. To improve robustness, recent approaches incorporate runtime monitoring of failures to trigger policy correction or retry [duan2024manipulateanythingautomatingrealworldrobots, etukuru2024robotutilitymodelsgeneral, dai2024racer]. In this work, we advance automatic failure detection by scaling structured training data and demonstrate improved integration with manipulation policies.
Robotic Failure Detection Methods. Early rule-based failure detection methods [de1998execution, gianni2011unified] struggle to generalize beyond predefined task structures. Learning-based approaches address this limitation and can be broadly categorized based on whether they require robot failure data for training.
Training-free methods. One line of work performs out-of-distribution (OOD) detection [canwedetectfailurewithoutfailuredata2025] or temporal inconsistency detection [agia2024unpackingfailuremodesgenerative] over internal policy representations to flag failures without failure supervision. Another line prompts LLMs or VLMs for failure assessment [ahmad2025unifiedframework] using techniques such as hierarchical CoT reasoning [liu2023reflect, agia2024unpackingfailuremodesgenerative] or constraint-aware visual programming [zhou2024code]. While these approaches avoid collecting robot failure data, they rely on policy-specific signals or prompt engineering, and cannot learn from robotic failure data for better accuracy.
Training-based methods. Recent works [duan2025aha, ifailsense2026, nvidia2026cosmosreason2, gu2025safe, pmlr-v232-du23b, armor2026] fine-tune VLMs as failure detectors using annotated trajectories, but mostly target execution-stage verification. SuccessVQA [pmlr-v232-du23b] performs coarse task-level success assessment. AHA [duan2025aha] and I-Fail-Sense [ifailsense2026] compress multi-view inputs into a single concatenated image for detecting subtask-level execution failures or instruction–behavior misalignment. SAFE [gu2025safe] relies on internal policy representation to train a failure classifier. Cosmos-Reason [nvidia2026cosmosreason2] addresses general embodied reasoning trained with supervised fine-tuning and reinforcement learning, where failure detection is one downstream task. ARMOR [armor2026] uses multi-round self-refinement with separate detection and reasoning heads, but it trains only on post-execution failures and requires multiple inference passes per sample. Compared to prior methods, our Guardian model leverages large-scale failure reasoning data to enable multi-view, explicit reasoning for unified planning and execution verification, achieving state-of-the-art performance.
Robot Failure Datasets. Collecting real-world robot failures at scale is challenging: policy rollouts are time-consuming, potentially unsafe, and require extensive manual annotation. RoboFail [liu2023reflect] provides a hand-crafted dataset spanning simulation and real settings, but covers limited tasks and failure modes. ViFailback [zeng2025vifailback] focuses on single-view, single-embodiment real-world diagnosis and requires substantial teleoperation effort. To reduce collection cost, several works rely on synthetic failure generation. Sentinel [agia2024unpackingfailuremodesgenerative] induces failures via out-of-distribution rollouts but covers only four tasks. SAFE [gu2025safe] utilizes the final sparse reward for policy rollouts and lacks dense failure supervision for each step. AHA [duan2025aha] perturbs trajectories in RLBench [james2019rlbenchrobotlearningbenchmark], generating large-scale purely simulated data, yet excludes high-level planning failures, and the dataset has not been publicly released. RoboFAC [robofac2025] adds reasoning annotations in simulation, with only a limited real-world subset via manual teleoperation. I-Fail-Sense [ifailsense2026] synthesizes failures from RLBench and DROID, but focuses primarily on semantic mismatches, leaving planning and low-level control errors underexplored. In contrast, we propose an automated pipeline that generates diverse planning and execution failures across both simulation and real robots, producing realistic failure modes at scale with multi-view observations and fine-grained, step-by-step reasoning supervision.
III FailCoT: Cross-Environment Robot Failure Reasoning Datasets
III-A Data Sources
Simulated data enables controlled failure generation through procedural perturbations [duan2025aha], while real robot data reduces the sim-to-real gap but requires substantial human supervision [liu2023reflect]. To balance precise control and real-world fidelity, we use both simulated and real-robot datasets to construct robot failure datasets. We propose an automated method that derives planning and execution failures directly from successful demonstrations, avoiding manual failure collection. In both domains, tasks are decomposed into subtasks with corresponding video segments, which form the basis for generating failures. Fig. 1 (middle row) illustrates successful episodes from the simulated and real robot datasets.
Simulated Data. We use the RLBench [james2019rlbenchrobotlearningbenchmark] simulator, selecting 52 tasks from RLBench-18Task [shridhar2022perceiveractormultitasktransformerrobotic] and GemBench [garcia2025generalizablevisionlanguageroboticmanipulation] benchmarks in our training data. For each task, we generate successful scripted trajectories with varied object placements and segment them into subtasks following [garcia2025generalizablevisionlanguageroboticmanipulation].
Real Robot Data. We use BridgeDataV2 [walke2024bridgedatav2datasetrobot] with ECoT annotations  [zawalski2025roboticcontrolembodiedchainofthought], which provide fine-grained subtasks and object labels using large VLMs. We further clean these annotations automatically using heuristics and Mistral-Small-3.1-24B [mistralsmall] to filter episodes with missing targets or unreliable bounding boxes. To increase the number of successful trajectories, we augment data by reversing successful executions when applicable, by swapping their start and end images, and updating the associated instructions accordingly (e.g., “open drawer” becomes “close drawer”, “flip pot upright” becomes “flip pot upside down”). This yields approximately 20% additional successful demonstrations.
III-B Automated Failure Data Generation
We design failure modes based on established failure taxonomies [liu2023reflect, duan2025aha] and analysis of robot policy failures [wu2025robomind]. The failures are categorized into two types: planning and execution. A planning error denotes an incorrect decomposition of a task into subplans, whereas an execution error reflects unsuccessful completion of a subplan.
Planning Failures. As shown in Fig. 1 (top row), we construct five types of planning failures:
  1. (1)
    Wrong object manipulated – some subtasks manipulate the wrong object.
  2. (2)
    Wrong object state or placement – some subtasks select the wrong target location, or state for the correct object.
  3. (3)
    Wrong order – one or several subtasks are not in the correct order, violating causal dependencies.
  4. (4)
    Missing subtask – required subtasks are missing from the plan, breaking task completeness.
  5. (5)
    Contradictory subtasks – some subtasks conflict with each other.
Types 1-3 are generated using an LLM (Mistral-Small-24B) to subtly alter the plan, while types 4-5 are created through rule-based perturbations. Each planning example comprises the task instruction, plan, and the initial front-view image.
Execution Failures. In simulation, we directly perturb subtask-level actions (Fig. 1, bottom left), leveraging the simulator’s precise control. A randomly selected subtask on the trajectory is modified using four failure modes:
  1. (1)
    No gripper close – the gripper is correctly positioned to grasp the object, but it fails to close its jaws.
  2. (2)
    Wrong object state or placement – the correct object is manipulated but ends in an incorrect state or placement.
  3. (3)
    Wrong object manipulated – the wrong object is used.
  4. (4)
    Imprecise grasping/pushing – the gripper attempts to grasp or push the correct object by moving toward it and closing its jaws, but misses it due to inaccurate positioning.
For real robot data, modifying actions directly is impractical due to current limitations of image editing and generation models. Therefore, we perturb the subtask text instruction paired with the pre-recorded trajectory segment (Fig. 1, bottom right) without direct robot control:
  1. (1)
    Task-execution semantic mismatch — an LLM (prompted with the original instruction and visible objects), or a rule-based preposition swap, generates a semantically altered instruction while preserving the start/end images.
  2. (2)
    Revert action — keep the instruction unchanged; replace the end image with the start one to show no progress.
Each execution example contains the task and subtask descriptions, plus pre-/post-action multi-view images.
Figure 2: Planning and execution examples with CoT.
III-C Chain-of-Thought (CoT) Generation
CoT reasoning has shown promise in improving the interpretability and performance of VLMs [zhang2024improve]. Therefore, we further explore whether reasoning can help failure detection. We introduce an automatic method to generate step-by-step CoTs for training reasoning models. For each sample, we first collect the object category, spatial location, and robot state from the RLBench simulator or from ECoT [zawalski2025roboticcontrolembodiedchainofthought] annotations, together with the corresponding failure reason. We then prompt a large reasoning-capable VLM (InternVL3-38B) [zhu2025internvl3exploringadvancedtraining] to generate step-by-step reasoning traces based on the initial text–image inputs and the aforementioned information. For planning samples, the model is instructed to sequentially verify each subtask and subsequently analyze the overall plan. For execution samples, the model is guided to describe the pre- and post-action images before assessing subtask completion. The reasoning trace contains 118 tokens on average. Fig. 2 illustrates training examples with chain-of-thoughts verifying plan correctness and subtask completion.
III-D Real-Robot, Policy-Driven Data Collection
To further support realistic evaluation, we curate UR5-Fail, a real-robot dataset, collected using a UR5 arm with three cameras. We run the 3D-LOTUS++ policy [garcia2025generalizablevisionlanguageroboticmanipulation] on 16 unique tasks, recording initial and final multi-view images for each subtask. Subtasks are manually labeled as success or failure to obtain execution failure data. For planning failures, we annotate ground-truth plans and generate failures using the method described in Sec. III-B. Unlike RoboFail [liu2023reflect], which is single-view and relies solely on teleoperation, UR5-Fail is three-view and features autonomous policy rollouts yielding more realistic failures.
III-E Dataset Statistics and Evaluation
FailCoT (RLBench-Fail, BridgeDataV2-Fail) and UR5-Fail, contain balanced success/failure examples across both planning and execution, with reasoning traces. FailCoT is split into training, validation, and test sets, with the validation and test sets featuring unseen tasks/environments to evaluate generalization, see Table I top. Table I bottom compares UR5-Fail with two existing real-world datasets and shows a more balanced distribution between execution and planning.
To measure the quality and diversity of our synthetic datasets, i.e., whether the generated failures reflect real policy execution, we run the 3D-LOTUS++ policy [garcia2025generalizablevisionlanguageroboticmanipulation] on 92 RLBench tasks and manually annotate failure modes for 3 failure episodes per task. As shown in Fig. 3, our designed failure modes reflect real failures, and the overall distribution of our synthetic and real failures remains similar.
TABLE I: Statistics of the FailCoT dataset and three real-world robot failure detection benchmarks. Our constructed datasets (FailCoT and UR5-Fail) contain balanced success and failure cases covering both execution and planning errors.
Training Validation Test
Dataset Env. Exec Plan Exec Plan Exec Plan
FailCoT (Ours)
   RLBench-Fail Sim 12358 5808 1000 500 1000 500
   BridgeDataV2-Fail Real 7830 4880 1000 500 1000 500
Real-World Robot Failure Detection Benchmarks
UR5-Fail (Ours) Real - - - - 140 140
RoboFail [liu2023reflect] Real - - - - 153 30
RoboVQA [sermanet2024] Real - - - - 357 -
Figure 3: Failure mode distributions in real executions and our constructed data.
IV Guardian: A multi-view reasoning VLM for robot failure detection
Figure 4: Left: Overview of the Guardian model architecture. Right: Integration of Guardian model into a robot manipulation pipeline for planning and execution verification. The failure reasoning trace is reinjected to help with replanning.
IV-A Problem Formulation
We formulate robot failure detection as a visual question answering problem. For planning verification, given a high-level task instruction $T$, a proposed plan $P=(P_{1},\cdots,P_{N})$, and the initial visual context $I_{\text{start}}$, the model $\text{VLM}_{\text{plan}}$ must decide whether the plan is correct or not:
$\text{VLM}_{\text{plan}}(I_{\text{start}},T,P)\rightarrow B_{\text{plan}}$ (1)
where $B_{\text{plan}}\in\{0,1\}$ indicates planning success.
For execution verification, given the task goal $T$, a subtask description $P_{i}$, and the visual observations before and after execution, $I_{\text{start}}$ and $I_{\text{end}}$, the model $\text{VLM}_{\text{exec}}$ similarly outputs
$\text{VLM}_{\text{exec}}(I_{\text{start}},I_{\text{end}},T,P_{i})\rightarrow B_{\text{exec}}$ (2)
where $B_{\text{exec}}\in\{0,1\}$ indicates execution success.
IV-B Model Architecture
The Guardian model is built upon the open-source VLM InternVL3-8B [zhu2025internvl3exploringadvancedtraining]. As shown in Fig. 4 (left), it comprises three components: a text tokenizer that converts text into discrete token embeddings, a visual encoder (InternViT-300M) that transforms individual images into visual embeddings, and a transformer-based LLM (Qwen2.5-7B) that processes the concatenated multimodal tokens to predict the answer.
Rather than concatenating multiple images into a single grid-based image as in prior work [duan2025aha, ifailsense2026], Guardian processes each image independently through the visual encoder. This design preserves fine-grained spatial details within each image and allows to explicitly reason about spatial and temporal changes for more accurate failure detection. Furthermore, instead of directly outputting classifications [duan2025aha, pmlr-v232-du23b, ifailsense2026], Guardian generates an explicit reasoning trace before concluding success or failure.
IV-C Model Training
We fine-tune Guardian on FailCoT using parameter-efficient Low-Rank Adaptation (LoRA) [hu2022lora], while freezing the visual encoder. Training minimizes cross-entropy loss for next-token prediction.
Although CoT has shown promise to improve performance, it brings additional computation overhead. Inspired by prior work [chen2025trainingstrategiesefficientembodied], we explore three strategies for incorporating CoT into failure detection: (1) Vanilla: a baseline model trained and evaluated to directly predict final answers (A) without CoT; (2) Thinking: the model is trained and inferred with explicit reasoning, always generating CoT before A; (3) Dropout: in training, the model alternates between generating CoT+A and directly predicting A, while at test time, only A is produced. Results in Sec. V-C show that adding reasoning traces consistently improves performance. The Thinking strategy performs best but increases inference time, while Dropout offers a better speed-accuracy trade-off.
IV-D Integration into Robotic Manipulation Framework
Guardian can be seamlessly plugged into existing robotic manipulation pipelines as a verification layer without requiring any architectural modification. Without loss of generality, consider a modular robotic manipulation framework. As shown in Fig. 4 (right), Guardian can be inserted at each planning and subtask execution step to detect potential failures. Upon detection, it can trigger replanning or re-execute the corresponding motion policy to facilitate recovery, and use its fine-grained failure reasoning as a hint to better replan.
V Experiments
V-A Experimental Setup
Evaluation datasets. Our main evaluation focuses on three unseen real-world benchmarks: RoboFail [liu2023reflect], UR5-Fail, and RoboVQA [sermanet2024]. RoboFail is a manually curated single-view UR5 failure dataset. UR5-Fail is our constructed multi-view real-robot dataset. RoboVQA (RVQA) is single-view and spans three embodiments: an Everyday Robots mobile manipulator, a human arm, and a human using a grasping tool. [^1] We use the RoboVQA test split restricted to execution success prediction. The original dataset also contains “planning” questions, but these focus on next-action/state prediction rather than plan verification. In ablations, we additionally report results on FailCoT testing splits. We use average classification accuracy as the metric.
Implementation details. We fine-tune models using LoRA (rank 16, effective batch size 16) with AdamW (weight decay 0.05), bf16 precision, and a cosine schedule peaking at $4\times 10^{-5}$. Training is conducted on H100 GPUs using FailCoT unless otherwise specified. For RLBench-Fail, we randomly sample one or four views during training to mitigate view-specific overfitting. The best checkpoint is selected via validation accuracy.
TABLE II: Comparison of failure detection models on unseen real-world benchmarks. Execution and planning accuracies are reported. denotes numbers from the original paper.
Model Trained on RoboFail [liu2023reflect] UR5-Fail RVQA [sermanet2024]
FailCoT Exec Plan Exec Plan Exec
Closed-Source VLM
GPT-4o 0.80 0.67 0.77 0.85 0.79
GPT-4o +Sentinel-Video-QA [agia2024unpackingfailuremodesgenerative] 0.80 0.63 0.76 0.62 0.66
Robotic Failure Detection VLMs
RoboFAC-7B [robofac2025] 0.25 0.05 0.54 0.02 0.52
AHA-13B [duan2025aha] 0.64 - - - -
I-Fail-Sense-3B [ifailsense2026] 0.43 0.67 0.47 0.46 0.53
Cosmos-Reason2-8B [nvidia2026cosmosreason2] 0.78 0.53 0.59 0.67 0.76
CLIP+MLP [pmlr-v139-radford21a] 0.42 0.43 0.51 0.51 0.52
I-Fail-Sense-3B [ifailsense2026] 0.76 0.52 0.55 0.6 0.58
Cosmos-Reason2-8B [nvidia2026cosmosreason2] 0.82 0.70 0.65 0.83 0.77
Guardian-8B 0.86 0.70 0.77 0.89 0.85
V-B Comparison with State of the Art
Compared methods. We compare against GPT-4o [openai2024gpt4ocard] and specialized robotic failure detectors including Cosmos-Reason2-8B [nvidia2026cosmosreason2], AHA-13B [duan2025aha], RoboFAC-7B [robofac2025], I-Fail-Sense-3B [ifailsense2026], Sentinel-Video-QA [agia2024unpackingfailuremodesgenerative], and CLIP+MLP [pmlr-v139-radford21a]. AHA results come from the original paper, as the model is not publicly released. For the other methods, we run their released checkpoints or train models with the released codebase.
Results. Table II reports performance on the test sets. GPT-4o achieves strong performance due to its scale and general reasoning ability. However, applying the Sentinel-Video-QA [agia2024unpackingfailuremodesgenerative] self-interrogation prompting degrades accuracy as it constrains the reasoning ability of the original model. Models trained exclusively on simulated failures (RoboFAC [robofac2025] and AHA [duan2025aha]) show limited transfer to real-robot benchmarks. Both rely on simulation-only perturbations, which likely restrict generalization to unseen real-world manipulators and sensor noise. I-Fail-Sense [ifailsense2026] is trained on both simulation and real-world trajectories, but its supervision focuses primarily on semantic misalignment detection rather than structured planning or low-level control failures, which likely limits its performance on the benchmarks. Cosmos-Reason [nvidia2026cosmosreason2] performs competitively, reflecting strong physical reasoning capabilities, but it is optimized for broad embodied reasoning rather than only for failure verification.
Since prior failure detection models are trained on different datasets, we further fine-tune representative open-source models on the same FailCoT dataset to isolate the effects of data and architecture (Table II). We also include a lightweight CLIP+MLP baseline, which performs substantially worse, highlighting the necessity of large vision–language models. Notably, training on FailCoT consistently improves all methods, underscoring the importance of well-curated, cross-environment data with broad failure coverage.
Guardian achieves the strongest overall performance across RoboFail, UR5-Fail, and RoboVQA. Compared to I-Fail-Sense [ifailsense2026], Guardian preserves multi-view spatial structure and produces explicit chain-of-thought reasoning, enabling structured subtask-level verification. Compared to Cosmos-Reason, which is pretrained for embodied reasoning and primarily developed in single-view settings, Guardian leverages an InternVL backbone with explicit multi-view supervision, which likely explains its stronger fine-grained failure detection performance even under identical training data.
TABLE III: Impact of the Guardian fine-tuning data mix on the binary accuracy averaged over planning and execution.
Training Data RLBench BDV2 Robo UR5 Robo RLBench BDV2 -Fail -Fail -Fail -Fail -VQA 0.65 0.69 0.65 0.73 0.75 0.82 0.70 0.69 0.72 0.66 0.65 0.86 0.71 0.68 0.77 0.85 0.88 0.78 0.83 0.85
Figure 5: FailCoT training data size impact on the binary accuracy averaged over planning and execution.
V-C Failure Data Ablations
We next analyze how training data composition (simulation and real data), and dataset scale influence cross-environment generalization while keeping the architecture fixed.
Data composition. Table III compares InternVL3-8B without fine-tuning, with single-source training, and with the full FailCoT dataset. Without fine-tuning, performance is moderate. Training only on RLBench-Fail improves simulated results but transfers weakly to real-robot datasets. Training only on BridgeDataV2-Fail (BDV2) improves real-world performance but shows limited simulation transfer. Combining both datasets yields consistent gains across RoboFail, UR5-Fail, and RoboVQA. The same trend holds for other architectures in Table II, emphasizing the importance of cross-environment composition and broad coverage of failure supervision.
Dataset scaling. Fig. 5 demonstrates consistent scaling behavior as the amount of FailCoT data increases. Performance on in-domain and unseen real-world datasets improves steadily with dataset size. This indicates that scaling structured failure generation remains a promising direction for improving cross-environment generalization.
TABLE IV: Comparison of training and test-time strategies with and without CoT (Sec. IV-C). “A” denotes the final answer. We report the binary accuracy and inference time (seconds/sample) on one H100 averaged across all test sets.